OpenStudy (anonymous):

the population was at 3046160 and had an increase of 1304226 now its asking me for the percent in change how can i figure this out?

OpenStudy (nikato):

Increase divided by original

OpenStudy (nikato):

Oh, and then multiply by 100

hero (hero):

First add the population increase to the old population to determine the new population: \[3046160 + 1304226\]

hero (hero):

Then use the formula for Percent Change: \[\frac{\% \Delta}{100} = \frac{|a - b|}{a}\] Where a = old amount = 3046160 b = new amount = 3046160 + 1304226

OpenStudy (nikato):

Y make it more complicated? @Hero

hero (hero):

I'm not. That's how you solve it

hero (hero):

How do you propose that it be solved?

hero (hero):

To me, it's not complicated at all. The explanation only seems complicated. But it isn't at all.

hero (hero):

You're probably just unfamiliar with some of the mathematical symbols

OpenStudy (anonymous):

uhm well i tried both formulas how about 57 % ?

OpenStudy (nikato):

Not quite. Maybe u should check ur calculations again. I got 42.82 or u can round to 43%

hero (hero):

@nikato is correct. We both got approximately 43% Do you mind showing your work @suzy4321 so that we can see what you did wrong?

OpenStudy (anonymous):

yeah i subtracted a-b and then divided by a 3064160-1304226 / 3064160=0.57

hero (hero):

I see.... Well, you misinterpreted the value of b b = 3046160 + 1304226

hero (hero):

Also you need to compute |a - b| not just a - b

OpenStudy (anonymous):

i see how u got the new answer and thanks for clarifying your both correct